OpenStudy (anonymous):

1. A pair of congruent angles are described as follows. The measure of one angle is 20 more than twice a number, and the other angle is 68 less than 4 times the same number. What is the size of the two angles? a. Define the variable: (1 point) b. Find an expression for the first and second angle. (1 point) c. Write the equation that represents this situation. (2 points) d. Solve for x. (2 points) e. Answer in complete sentence: (1 point) PLZZ I need help I just don't understand this question ill fan

Directrix (directrix):

Let x be the number. Let <1 and <2 be the angles m<1 = 2x + 20 m<2 = 4x -68 If the two angles are congruent, then m<1 = m<2. Solve for x: 2x + 20 = 4x - 68 Post what you get, okay?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@directrix I really cant figure it out

OpenStudy (anonymous):

@goprogo501 this is the one

Directrix (directrix):

You can do this: 2x + 20 = 4x - 68 Solve for x 2x - 4x = -68 - 20 ?

OpenStudy (anonymous):

oh I didn't know I had to do that

Directrix (directrix):

What is 2x - 4x =

OpenStudy (anonymous):

-2x

Directrix (directrix):

What is -68 - 20 =

OpenStudy (anonymous):

-88

Directrix (directrix):

2x + 20 = 4x - 68 2x - 4x = -68 - 20 -2x = -88 x = ?

OpenStudy (anonymous):

-44

Directrix (directrix):

-88 divided by -2 = + 44

Directrix (directrix):

x = 44

Directrix (directrix):

Now, you have all the parts you need to complete the posted question.

OpenStudy (anonymous):

oh ok THANK YOU I appreciate it a lot :D, and I think I understand how you did it too

Directrix (directrix):

You are welcome.